OpenStudy (anonymous):

5. Which equation can be used to solve the following word problem: Steve has 5 more nickels than dimes, and the total value of his coins is $2.05. How many nickels does he have? Let n represent the number of nickels. a. 0.10(n – 5) + 0.05n = 2.05 b. 0.10(n – 5) + 0.05n = 2.05 c. 0.10n + 0.05(n – 5) = 2.05 d. 0.10n + 0.05n + 5 = 2.05

OpenStudy (anonymous):

@arabpride I need help again llol

OpenStudy (perl):

we are given that n = 5+ d , where n = # of nickels, d = # of dimes. and we are given that .05*n + .10*d = 2.05

OpenStudy (perl):

after substituting you should get a)

OpenStudy (perl):

your a) and b) look exactly alike

OpenStudy (anonymous):

im, confused?